# Quantum Feature Space of a Qubit Coupled to an Arbitrary Bath

### v3

@@ Abstract
Qubit control protocols have traditionally leveraged a characterisation of the qubit-bath coupling via its power spectral density. Previous work proposed the inference of noise operators that characterise the influence of a classical bath using a grey-box approach that combines deep neural networks with physics-encoded layers. This overall structure is complex and poses challenges in scaling and real-time operations. Here, we show that no expensive neural networks are needed and that this noise operator description admits an efficient parameterisation. We refer to the resulting parameter space as the quantum feature space of the qubit dynamics resulting from the coupled bath. We show that the Euclidean distance defined over the quantum feature space provides an effective method for classifying noise processes in the presence of a given set of controls. Using the quantum feature space as the input space for a simple machine learning algorithm (random forest, in this case), we demonstrate that it can effectively classify the stationarity and the broad class of noise processes perturbing a qubit. Finally, we explore how control pulse parameters map to the quantum feature space.
Keywords: quantum control, quantum features, classification, clustering
@@ 1 Introduction
Noise in quantum computers that results in qubits losing coherence (decoherence) is due to unwanted bath coupling [1]. One way to mitigate decoherence is to decouple the qubit from the bath by applying a control field as a series of pulses [2]. This technique was first studied in nuclear magnetic resonance imaging [3, 4]. Within this framework, control pulses are applied for a time $\tau$ and then turned off for a time $\tau$, and this pattern is repeated $N$ times. This oscillating pulse sequence design is the Carr-Purcell-Meiboom-Gill (CPMG) sequence [5, 6].
@@
Álvarez and Suter (AS) proposed a modified CPMG sequence to perform qubit spectroscopy [7]. AS pulses are part of a family of pulses used for dynamical decoupling noise spectroscopy, and they can be used to infer qubit coherence as a function of time. These empirical coherence data, obtained by AS pulses, are fitted to theoretical coherence curves corresponding to parametric noise spectra [8, 9]. However, this method requires an accurate model of system dynamics [10], which can be challenging to develop in general [11]. Furthermore, if an inappropriate parametric noise spectrum is chosen a priori, the estimated noise spectrum will correspond to an inaccurate model of system dynamics, and subsequently, control pulses optimised using this system model will likely be suboptimal [11].
@@
More direct methods of noise spectroscopy include techniques such as two-point correlation, which measures the correlation function of the noise by implementing gates at different times and measuring the change in the qubit state [12]. The Fourier transform of this correlation function gives the noise spectrum. However, this method can be resource intensive, time-consuming, computationally expensive, and requires that the noise be stationary [8, 10, 13, 14].
@@
The authors of [14] developed a quantum noise spectroscopy method that used the Fourier transform of the measurements of the free-induction decay of the coherence curves. The technique accurately recovered the correct noise spectra and outperformed previous decoupling schemes while significantly reducing experimental overhead. However, the method makes several assumptions that limit the applicability of the work. For example, the authors assume the qubit is only subject to dephasing, where the qubit thermal relaxation process occurs over a much longer time scale than phase randomisation, and that the frequency fluctuations of a qubit are subject to stationary zero-mean Gaussian noise.
@@
In [10], researchers instead used deep learning techniques to infer accurate qubit noise spectra. The method used coherence curves as input to a recurrent neural network with a feedforward neural network head to estimate the noise spectrum. They showed that the model could accurately estimate the qubit noise spectrum using only the coherence curves and outperformed the standard noise spectroscopy techniques. Researchers in [13] further demonstrated the effectiveness of deep learning by using it to reconstruct the power spectral density (PSD) of an ensemble of carbon impurities around a diamond’s nitrogen-vacancy centre.
@@
Youssry et al. [15] introduced a grey-box model separating control and system-bath dynamics. In this work, time-dependent Hamiltonians and unitaries specified the control dynamics, and a form of recurrent neural network (a gated recurrent unit) was used to predict the system-bath dynamics associated with control pulses. The input to this neural network was parameterised control pulses as a sequence of vectors. The model output was the parameterisation of a noise operator. The authors showed that the grey-box model could be used for qubit control pulse optimisation and qubit noise spectroscopy. This grey-box model approach was later used for qubit control in the presence of a non-Markovian bath [16]. The formalism used in this work is discussed in more detail in Sec. 2.
@@
Despite the success of [10, 13, 15, 16], these approaches are limited by their use of deep learning models, specifically the expense of training deep learning models and the need for large datasets. Expensive models are undesirable as noise processes in quantum systems can drift over time [17, 18], necessitating new training data and model retraining potentially daily or even hourly. Additionally, noise profiles may change from one physical qubit to another [19], so a specific model may be required for each qubit, which does not scale. Furthermore, specific to the grey-box method of [15, 16], the white-box aspect (the so-called physics-informed aspect of the network) scales exponentially with the dimensions of the quantum system, adding to the challenge of moving beyond toy problems.
@@
We introduce in this paper quantum feature spaces, which encode the bath’s influence on system dynamics while making minimal assumptions about the bath such that it can be used beyond toy problems and have utility for a broad range of physically realised qubits. The quantum feature space is deduced from observable expectations using efficient linear regression models. We then show that no expensive algorithms are needed to classify the noise affecting a qubit; instead, we use the Euclidean distance within the quantum feature space for classification. Fig. 1 demonstrates further uses for the quantum feature space. The feature space can be used as the input space to simple machine learning algorithms for noise classification, or the space can enable control pulse optimisation via visualising how control pulse parameters manifest themselves in the quantum feature space.
Figure 1: The diagram illustrates an iterative process for characterising and learning noise properties in quantum systems. Starting with a power spectral density (PSD) (top), noise realisations are generated (right) and combined with control pulses in a quantum simulation. The outcomes of these simulations are mapped onto a quantum feature space (bottom), where data points encode the effects of the noise. This feature space is used by regression models such as decision trees to infer key noise properties, including type and stationarity (left). These inferred properties refine the PSD model used in subsequent iterations. This closed-loop framework enables the classification and clustering of noise processes. It facilitates tailored noise mitigation strategies, leveraging machine learning techniques like decision trees, k-means clustering, and neural networks for enhanced quantum control.
@@ 2.1 Open System Dynamics
The Hamiltonian describing a controlled qubit (the system) interacting with a bath is,
$H(t)=H_{ctrl}(t)+H_{SB}(t)$ (1)
where the control Hamiltonian is,
@@
where $\Omega$ is the qubit’s energy gap and $f_{j}(t)$ is a control pulse along the given axis, and the system-bath Hamiltonian is,
$H_{SB}(t)=\sum_{j=x,y,z}\sigma_{j}\otimes B_{j}(t)$ (3)
with $B_{j}(t)=\hat{B}_{j}(t)+\beta_{j}(t)I_{B}$ capturing a quantum bath via the operator $\hat{B}_{j}(t)$ and classical noise via the stochastic process $\beta_{j}(t)$.
@@
We are interested in the expected value of a system (qubit) observable $O$ at time $T$ given an initial system (qubit) state $\rho$, where $T$ represents the total time of the control pulse sequence. This is given by,
@@
$\mathbb{E}\{O(T)\}_{\rho}=\left\langle\mathrm{Tr}\left[U(T)\left(\rho\otimes\rho_{B}\right)U(T)^{\dagger}O\right]\right\rangle$ (4)
[…]
We then decompose the total system-bath evolution into a product of system-bath dynamics and control dynamics. To do this, Paz-Silva, Norris, and Lorenza [20] used a so-called toggling frame, defined by the time-ordered control unitary,
[…]
The system-bath Hamiltonian in this toggling frame is,
$H_{\mathrm{SB}}^{\mathrm{tog}}(t)=U_{\mathrm{ctrl}}^{\dagger}(t)H_{\mathrm{SB}}(t)U_{\mathrm{ctrl}}(t)$ (6)
with the associated unitary,
[…]
It can be shown that,
${U}_{\mathrm{SB}}(t)=U_{\mathrm{ctrl}}(t)U_{\mathrm{SB}}^{\mathrm{tog}}(t)U_{\mathrm{ctrl}}^{\dagger}(t)$ (8)
which allows us to rewrite Eq. 4 as,
@@
$\mathbb{E}\{O(T)\}_{\rho}=\mathrm{Tr}\left[V_{O}(T)U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}(T)^{\dagger}O\right]$ (9)
where
@@
$V_{O}(T)=\left\langle O^{-1}{U}_{\mathrm{SB}}^{\dagger}(T)O{U}_{\mathrm{SB}}(T)\right\rangle_{B}$ (10)
with $\left\langle\cdot\right\rangle_{B}=\left\langle\mathrm{Tr}_{B}\left[\cdot\rho_{B}\right]\right\rangle$ representing classical averaging over the partial trace taken with respect to the bath. Physically, we interpret $V_{O}$ as a so-called noise operator that encodes the influence of the system-bath interaction on qubit dynamics.
2.2 Decomposition of the System-Bath Operator
When our system-bath Hamiltonian consists of only classical noise processes, one can move $O^{-1}$ in Eq. 10, and write,
$V_{O}=O^{-1}\left\langle U_{\mathrm{SB}}^{\dagger}OU_{\mathrm{SB}}\right\rangle$ (11)
where the time arguments $T$ have been omitted for brevity. Before continuing, it is important to note some of the mathematical properties of the $V_{O}$ operator as they subsequently inform restrictions on its parameterisation. Specifically, $\left\langle U_{\mathrm{SB}}^{\dagger}OU_{\mathrm{SB}}\right\rangle$ is trace bounded by one, is Hermitian, and has real eigenvalues between $-1$ and $1$ [15].
An efficient parameterisation of $V_{O}$ is provided by its eigendecomposition. This eigendecomposition contains a diagonal matrix $D_{O}$ whose entries are real numbers in the interval $[-1,1]$ and whose sum also lies in the interval $[-1,1]$, and a general unitary matrix $Q$. Thus Eq. 11 can be rewritten as
$V_{O}=O^{-1}Q_{O}D_{O}Q_{O}^{\dagger}$ (12)
In the case of a qubit, we start with a factorisation of a general $2\times 2$ unitary operator,
$Q_{O}=\left[\begin{array}[]{cc}e^{i\psi_{O}}&0\\ 0&e^{-i\psi_{O}}\end{array}\right]\left[\begin{array}[]{cc}\cos\theta_{O}&\sin\theta_{O}\\ -\sin\theta_{O}&\cos\theta_{O}\end{array}\right]\left[\begin{array}[]{cc}e^{i\Delta_{O}}&0\\ 0&e^{-i\Delta_{O}}\end{array}\right]$ (13)
where $\psi_{O},\theta_{O},\Delta_{O}\in\mathbb{R}$ and the diagonal matrix $D_{O}$ of Eq. 12 is constructed as
$D_{O}=\left[\begin{array}[]{cc}\mu_{O}&0\\ 0&-\mu_{O}\end{array}\right]$ (14)
where $\mu_{O}\in[0,1]$. The matrix $D_{O}$ can be generalised to non-unital channels [15], but this paper will focus on unital channels, for which $\Phi(I)=I$ [21]. The restriction to unital channels in this study is motivated by the fact that for many physical qubit implementations, the decoherence timescale ($T2$) is observed to be significantly shorter than the energy relaxation timescale ($T1$[21, 22]. Unital channels effectively model this rapid decoherence, and are often used to describe the dynamics of qubits in the presence of a classical bath [21].
@@ 2.3 Deriving the Quantum Feature Space
The previous grey-box approach to qubit control of [15, 16] used expectations and control pulse amplitudes to train a model, both classical feature spaces. This motivated us to develop a method to deduce the $Q_{O}D_{O}Q_{O}^{\dagger}$ operators and their parameters from observable expectations. This allowed us to understand the properties of the qubit bath over time. The feature space formed from the parameters of $Q_{O}D_{O}Q_{O}^{\dagger}$ was hypothesized to be a richer feature space than those used in the previous work. By creating a rich feature space, we could classify noise processes without needing expensive deep learning models.
Substituting Eq. 12 into Eq. 9 gives,
$\mathbb{E}\{O(T)\}_{\rho}=\mathrm{Tr}\left[Q_{O}D_{O}Q_{O}^{\dagger}U_{\mathrm{ctrl}}\rho U_{\mathrm{ctrl}}^{\dagger}\right]$ (15)
To calculate the mapping between expectations and $Q_{O}D_{O}Q_{O}^{\dagger}$, one must understand a priori the generalised structure of $Q_{O}D_{O}Q_{O}^{\dagger}$. Recalling the definition of $Q_{O}$, we find $Q_{O}D_{O}Q_{O}^{\dagger}$ is,
$\left[\begin{array}[]{cc}\mu_{O}\cos(2\theta_{O})&-e^{2i\psi_{O}}\mu_{O}\sin(2\theta_{O})\\ -e^{-2i\psi_{O}}\mu_{O}\sin(2\theta_{O})&-\mu_{O}\cos(2\theta_{O})\end{array}\right]=\left[\begin{array}[]{cc}\gamma_{O}&\alpha_{O}-\beta_{O}i\\ \alpha_{O}+\beta_{O}i&-\gamma_{O}\end{array}\right]$ (16)
with $-1\leq\alpha_{O},\beta_{O},\gamma_{O}\leq 1$.
Now $U_{\mathrm{ctrl}}\rho U_{\mathrm{ctrl}}^{\dagger}$ must be Hermitian, with trace one, such that,
@@
We can now calculate the generalised form of $\mathbb{E}\{O(T)\}_{\rho}$ in terms of the newly introduced parameters,
$\mathbb{E}\{O(T)\}_{\rho}=2b_{\rho}\alpha_{O}+2c_{\rho}\beta_{O}+(2a_{\rho}-1)\gamma_{O}$ (18)
For each observable, there are six initial states, so to find solutions $\alpha_{O},\beta_{O},\gamma_{O}$ for each observable, we need to solve an over-determined system of six equations with three unknowns,
@@
The section has also shown how the quantum simulator takes a control sequence and stochastic realisation defined for $t$ timesteps and produces quantum feature space parameters ($\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$) and observable expectations ($\mathbb{E}\{O(T)\}_{\rho}$). Properties within the quantum feature space (such as the Euclidean distance) can then be used to estimate the underlying noise model and its parameters. The quantum feature space can also form the input to a simple machine learning algorithm (such as a random forest) or a small neural network, both with nine inputs for the single-qubit case (three parameters for each of the three observables). The output of the machine learning algorithm can be used to classify properties of the noise process, such as its stationarity and type, where the algorithm is trained on labels from the known simulated noise processes. The trained model can then be applied to an unknown noise process from a physical quantum computer to estimate the noise properties and parameters.

### v4

@@ Abstract
Qubit control protocols have traditionally leveraged a characterisation of the qubit-bath coupling via its power spectral density. Previous work infered noise operators that characterise the influence of a classical bath using a grey-box approach that combines deep neural networks with physics-encoded layers. This overall structure is complex and poses challenges in scaling and real-time operations. Here, we show that no expensive neural networks are needed and that this noise operator description admits an efficient parameterisation. We refer to the resulting parameter space as the quantum feature space (QFS) of the qubit dynamics resulting from the coupled bath. We show that the Euclidean distance defined over the QFS provides an effective method for classifying noise processes in the presence of a given set of controls. Using the QFS as the input space for a simple machine learning algorithm, we demonstrate that it can effectively classify the stationarity and the broad class of noise processes perturbing a qubit. Finally, we explore how control pulse parameters map to the QFS.
keywords:
quantum control, quantum features, classification, clustering
@@ 1 Introduction
Noise in quantum computers that results in qubits losing coherence (decoherence) is due to unwanted bath coupling [bravyi2018correcting]. One way to mitigate decoherence is to decouple the qubit from the bath by applying a control field as a series of pulses [viola1999dynamical]. This technique was first studied in nuclear magnetic resonance imaging [degen2017quantum, boss2016one]. Within this framework, control pulses are applied for a time $\tau$ and then turned off for a time $\tau$, and this pattern is repeated $N$ times. This oscillating pulse sequence design is the Carr-Purcell-Meiboom-Gill (CPMG) sequence [carr1954effects, meiboom1958modified].
@@
Álvarez and Suter (AS) proposed a modified CPMG sequence to perform qubit spectroscopy [alvarez2011measuring]. AS pulses are part of a family of pulses used for dynamical decoupling noise spectroscopy, and they can be used to infer qubit coherence as a function of time. These empirical coherence data, obtained by AS pulses, are fitted to theoretical coherence curves corresponding to parametric noise spectra [von2020two, sun2022self]. However, this method requires an accurate model of system dynamics [wise2021using], which can be challenging to develop in general [martina2023machine]. Furthermore, if an inappropriate parametric noise spectrum is chosen a priori, the estimated noise spectrum will correspond to an inaccurate model of system dynamics, and subsequently, control pulses optimised using this system model will likely be suboptimal [martina2023machine].
@@
More direct methods of noise spectroscopy include techniques such as two-point correlation, which measures the correlation function of the noise by implementing gates at different times and measuring the change in the qubit state [baroni2022nuclear]. The Fourier transform of this correlation function gives the noise spectrum. However, this method can be resource-intensive, time-consuming, computationally expensive, and requires that the noise be stationary [von2020two, wise2021using, martina2023deep, vezvaee2024fourier].
@@
The authors of [vezvaee2024fourier] developed a quantum noise spectroscopy method that used the Fourier transform of the measurements of the free-induction decay of the coherence curves. The technique accurately recovered the correct noise spectra and outperformed previous decoupling schemes while significantly reducing experimental overhead. However, the method makes several assumptions that limit the applicability of the work. For example, the authors assume the qubit is only subject to dephasing, where the qubit thermal relaxation process occurs over a much longer time scale than phase randomisation, and that the frequency fluctuations of a qubit are subject to stationary zero-mean Gaussian noise.
@@
In [wise2021using], researchers instead used deep learning techniques to infer accurate qubit noise spectra. The method used coherence curves as input to a recurrent neural network with a feedforward neural network head to estimate the noise spectrum. They showed that the model could accurately estimate the qubit noise spectrum using only the coherence curves and outperformed the standard noise spectroscopy techniques. Researchers in [martina2023deep] further demonstrated the effectiveness of deep learning by using it to reconstruct the power spectral density (PSD) of an ensemble of carbon impurities around a diamond’s nitrogen-vacancy centre.
@@
Youssry et al. [youssry2020characterization] introduced a grey-box model separating control and system-bath dynamics. In this work, time-dependent Hamiltonians and unitaries specified the control dynamics, and a form of recurrent neural network (a gated recurrent unit) was used to predict the system-bath dynamics associated with control pulses. The input to this neural network was parameterised control pulses as a sequence of vectors. The model output was the parameterisation of a noise operator. The authors showed that the grey-box model could be used for qubit control pulse optimisation and qubit noise spectroscopy. This grey-box model approach was later used for qubit control in the presence of a non-Markovian bath [youssry2022multi]. The formalism used in this work is discussed in more detail in Sec. 2.
@@
Despite the success of [wise2021using, martina2023deep, youssry2020characterization, youssry2022multi], these approaches are limited by their use of deep learning models, specifically the expense of training deep learning models and the need for large datasets. Expensive models are undesirable as noise processes in quantum systems can drift over time [seedhouse2025wavelet, nakamura2024gate], necessitating new training data and model retraining potentially daily or even hourly. Additionally, noise profiles may change from one physical qubit to another [sung2019non], so a specific model may be required for each qubit, which does not scale. Furthermore, specific to the grey-box method of [youssry2020characterization, youssry2022multi], the white-box aspect (the so-called physics-informed aspect of the network) scales exponentially with the dimensions of the quantum system, adding to the challenge of moving beyond toy problems.
@@
We introduce in this paper the quantum feature space (QFS), which encodes the bath’s influence on system dynamics while making minimal assumptions about the bath such that it can be used beyond toy problems and have utility for a broad range of physically realised qubits. The QFS is deduced from observable expectations using efficient linear regression models. We then show that no expensive algorithms are needed to classify the noise affecting a qubit; instead, we use the Euclidean distance within the QFS for classification. Fig. 1 demonstrates further uses for the QFS. The feature space can be used as the input space to simple machine learning algorithms for noise classification, or the space can enable control pulse optimisation via visualising how control pulse parameters manifest themselves in the QFS.
Figure 1: The diagram illustrates an iterative process for characterising and learning noise properties in quantum systems. Starting with a PSD (top), noise realisations are generated (right) and combined with control pulses in a quantum simulation. The outcomes of these simulations are mapped onto a QFS (bottom), where data points encode the effects of the noise. This feature space is used by regression models such as decision trees to infer key noise properties, including type and stationarity (left). These inferred properties refine the PSD model used in subsequent iterations. This closed-loop framework enables the classification and clustering of noise processes. It facilitates tailored noise mitigation strategies, leveraging machine learning techniques like decision trees, k-means clustering, and neural networks for enhanced quantum control.
@@ 2.1 Open System Dynamics
We start by modeling a single qubit perturbed by classical noise processes $\beta_{j}(t)$} to represent system-bath interactions (using units where $\hbar=1$). The total Hamiltonian is then
$H(t)=H_{\mathrm{ctrl}}(t)+H_{\mathrm{noise}}(t)$ (1)
with
@@
where $\Omega$ is the energy gap of the qubit and $f_{j}(t)$ are externally applied control pulses; and
$H_{\mathrm{noise}}(t)=\sum_{j=\{x,y,z\}}\beta_{j}(t)\sigma_{j}$ (3)
where $\beta_{j}(t)$ are stochastic processes that represent the noise.
@@
We are interested in the expected value of a system (a qubit) observable $O$ at time $T$ given an initial state of the system $\rho$, where $T$ represents the total time of the control pulse sequence. This is given by,
$\mathbb{E}\{O(T)\}_{\rho}=\left\langle\mathrm{Tr}\left[U(T)\rho U(T)^{\dagger}O\right]\right\rangle$ (4)
where $U(T)=\mathcal{T}e^{-i\int_{0}^{T}H(s)\,ds}$ is the time-ordered exponential of the time-dependent Hamiltonian in Eq. 1, and $\langle\cdot\rangle$ denotes classical averaging over realisations of random processes $\beta_{j}(t)$.
We are interested in decomposing the total evolution into a product of system-bath dynamics and system dynamics. This decomposition allows us to separate the influence of the bath on the system, which is crucial for understanding how noise affects qubit control and dynamics. To do this, we follow the work of Paz-Silva, Norris, and Lorenza [paz2017multiqubit] and Youssry, Pas-Silva, and Ferrie [youssry2020characterization] and move to an interaction picture with respect to the control Hamiltonian.
We start then with the time-ordered control unitary,
$U_{\mathrm{ctrl}}(t)=\mathcal{T}e^{-i\int_{0}^{t}H_{\mathrm{ctrl}}(s)\,ds}$ (5)
We then define the interaction Hamiltonian as follows,
$H_{\mathrm{I}}(t)=U_{\mathrm{ctrl}}^{\dagger}(t)H_{\mathrm{noise}}(t)U_{\mathrm{ctrl}}(t)$ (6)
with the following associated unitary,
$U_{\mathrm{I}}(t)=\mathcal{T}e^{-i\int_{0}^{t}H_{\mathrm{I}}(s)\,ds}$ (7)
Using this definition of the interaction Hamiltonian allows us to rewrite the total unitary as a product of the interaction unitary and the control unitary without assuming that the control and noise Hamiltonians commute (see Sec. 7.1 for more details). Specifically, we can write the total unitary as,
$U(t)=U_{\mathrm{ctrl}}(t)U_{\mathrm{I}}(t)=U_{\mathrm{ctrl}}(t)U_{\mathrm{I}}(t)U_{\mathrm{ctrl}}^{\dagger}(t)U_{\mathrm{ctrl}}(t)=\tilde{U}_{\mathrm{I}}(t)U_{\mathrm{ctrl}}(t)$ (8)
where $\tilde{U}_{\mathrm{I}}(t)=U_{\mathrm{I}}(t)U_{\mathrm{ctrl}}^{\dagger}(t)U_{\mathrm{ctrl}}(t)$ is a modified interaction unitary defined so that we can rewrite the expectation in Eq. 4 as
@@
$\mathbb{E}\{O(T)\}_{\rho}=\left\langle\mathrm{Tr}\left[\tilde{U}_{\mathrm{I}}(T)U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}^{\dagger}(T)\tilde{U}_{\mathrm{I}}^{\dagger}(T)O\right]\right\rangle$ (9)
using the cyclic property of the trace, we have
@@
$\mathbb{E}\{O(T)\}_{\rho}=\left\langle\mathrm{Tr}\left[U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}^{\dagger}(T)\tilde{U}_{\mathrm{I}}^{\dagger}(T)O\tilde{U}_{\mathrm{I}}(T)\right]\right\rangle$ (10)
Bringing the trace inside the expectation and moving the constants out, we have
@@
$\mathbb{E}\{O(T)\}_{\rho}=\mathrm{Tr}\left[U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}^{\dagger}(T)\left\langle\tilde{U}_{\mathrm{I}}^{\dagger}(T)O\tilde{U}_{\mathrm{I}}(T)\right\rangle\right]$ (11)
where we define the QFS operator as
$\tilde{O}(T)=\left\langle\tilde{U}_{\mathrm{I}}^{\dagger}(T)O\tilde{U}_{\mathrm{I}}(T)\right\rangle$ (12)
giving us,
$\mathbb{E}\{O(T)\}_{\rho}=\mathrm{Tr}\left[U_{\mathrm{ctrl}}(T)\rho U_{\mathrm{ctrl}}^{\dagger}(T)\tilde{O}(T)\right]$ (13)
This shows that the measured expectation depends on the system’s evolution under control ($U_{\mathrm{ctrl}}(T)$) and the bath’s influence, captured entirely by $\tilde{O}(T)$. Notice, that at no point did we assume $H_{\mathrm{ctrl}}(t),H_{\mathrm{noise}}(t)]=0$; instead, we used the interaction picture to isolate bath effects, consistent with standard techniques in open quantum systems.
Note, continuing forward, we will omit the time-dependence notation for clarity, and it is understood that the parameters and operators are meant to be evaluated at the end of the control pulse sequence, at time $T$, if the time-dependence is not explicitly stated.
@@ 2.2 Deriving the Quantum Feature Space
The previous grey-box approach to qubit control of [youssry2020characterization, youssry2022multi] used expectations and control pulse amplitudes, both classical feature spaces. This motivated us to develop a method to deduce the $\tilde{O}$ operator and its parameters from observable expectations as it was hypothesised that the feature space formed from the parameters of $\tilde{O}$ would be a richer feature space than those used in previous work. By creating a rich feature space, we could classify noise processes without needing expensive deep learning models.
As such, it is important to note some of the mathematical properties of $\tilde{O}$ as they subsequently inform restrictions on its parameterisation. Specifically, $\tilde{O}$ has zero trace, is Hermitian, and has real eigenvalues between 1-1 and $1$ [youssry2020characterization], resulting in the following form,
$\tilde{O}=\left[\begin{array}[]{cc}\gamma_{O}&\alpha_{O}-\beta_{O}i\\ \alpha_{O}+\beta_{O}i&-\gamma_{O}\end{array}\right]$ (14)
with $1\leq\alpha_{O},\beta_{O},\gamma_{O}\leq 1$.
Now $U_{\mathrm{ctrl}}\rho U_{\mathrm{ctrl}}^{\dagger}$ must be Hermitian, with trace one, and positive semi-definite, giving us the following form,
@@
We can now calculate the generalised form of $\mathbb{E}\{O\}_{\rho}$ in terms of the newly introduced parameters,
$\mathbb{E}\{O\}_{\rho}=2b_{\rho}\alpha_{O}+2c_{\rho}\beta_{O}+(2a_{\rho}-1)\gamma_{O}$ (16)
For a single qubit (two-dimensional system), for each observable, we measure the observable expectation value for six initial states, specifically the up and down eigenstates for each of the three Pauli matrices ($\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$). This process yields an over-determined system of six linear equations with three unknowns, which we can solve to find the parameters $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$ for each observable $O$. The system is represented by the following equation:
@@
The section has also shown how the quantum simulator takes a control sequence and stochastic realisation defined for $T$ timesteps and produces QFS parameters ($\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$) and observable expectations ($\mathbb{E}\{O\}_{\rho}$). For a physical quantum computer, this section has also shown how to use observable expectations to deduce the QFS parameters using the over-determined system of equations in Eq. 17. Properties within the QFS (such as the Euclidean distance) can then be used to estimate the underlying noise model and its parameters. The QFS can also form the input to a simple machine learning algorithm (such as a random forest) or a small neural network, both with nine inputs for the single-qubit case (three parameters for each of the three observables). The output of the machine learning algorithm can be used to classify properties of the noise process, such as its stationarity and type, where the algorithm is trained on labels from the known simulated noise processes. The trained model can then be applied to an unknown noise process from a physical quantum computer to estimate the noise properties and parameters.